OpenStudy (anonymous):

(3x+8)^2 HELP PLEASE?????

OpenStudy (anonymous):

9x^2+48x+64

OpenStudy (anonymous):

a. 9x2 - 64 b. 9x2 + 64 c. 9x2 + 48x + 64 d. 9x2 - 48x + 64

OpenStudy (anonymous):

So C

OpenStudy (tyteen4a03):

@Lolovitz Please don't explain answers without explaining them. It doesn't help the student learn and is in fact against the OpenStudy Code of Conduct. (3x+8)^2 = (3x+8)(3x+8). FOIL.

OpenStudy (tyteen4a03):

@Kaggler How did you manage to get B?

OpenStudy (anonymous):

OHHHH

OpenStudy (anonymous):

@tyteen4a03 I did 3^2=9 = 9x^2 then 8^2=64 so I got 9x^2 + 64

OpenStudy (texaschic101):

do you understand the FOIL method ?

OpenStudy (anonymous):

But I get it now (: thank you guys!!!

OpenStudy (anonymous):

I used the formula (a+b)^2=a^2+2a*b+b^2

OpenStudy (tyteen4a03):

@kaggler That is not correct. Remember this identity in your heart: (a+b)^2 = a^2 + 2ab + b^2

OpenStudy (anonymous):

9x^2 + 48x +64? @tyteen4a03

OpenStudy (tyteen4a03):

@Kaggler Correct.

OpenStudy (anonymous):

Thank you!!!!!!!!! @tyteen4a03